# PoseScript: Linking 3D Human Poses and Natural Language

### v2

@@ A-B List of posecodes
Posecodes statistics. In Figures 16, 17, 18, 19, 20, 21 and 22 we show posecode statistics obtained over PoseScript-A. Specifically, circle areas represent the proportion of poses satisfying the corresponding posecode categorization for the associated keypoints. We use the black and grey colors to denote categorizations that are ignored in the captioning process. A black circle area means that the corresponding pose configuration is too ambiguous (e.g. when the relative distance between two body parts is close to 0, making the detection of the body parts’ relative position less obvious.). Grey circle areas denote trivial pose configurations (e.g. when a left body part is at the left of the associated right body part: this is the case most of the time). They correspond to posecode categorizations that apply to at least 60% of the poses. In contrast, posecode categorizations that describe less than 6% of the poses are defined as unskippable (i.e., such pose information cannot be randomly discarded during the posecode selection process), and are colored in orange. All other available posecodes categorizations, in blue, are skippable (i.e., such pose information can be randomly discarded during the posecode selection process). Equivalent information for super-posecodes is provided in Table IX.
@@ Appendix B More about the automatic captioning pipeline
Rules were mined based on PoseScript-A. We further reviewed all mined rules manually, to keep only the most meaningful and dispose of the following:
@@
Rotation-related information. As it uses 3D coordinates as input, the automatic captioning pipeline describes relative body part positioning, but does not inform about limb orientation or limb twist. Yet, this could be easily taken into account by designing new posecodes which computation would rely on joint angles (instead of joint coordinates). This is mostly interesting and worth describing for end limbs like lower arms (eg. to tell whether the inside of the lower arm is facing down or facing up, when the wrist is extended forward). Usually, twists of other limbs either induce visual twists of the end limbs connected to them, or result in different positions of other 3D keypoints. As such, in most of the case, the captioning of these 3D keypoints does implicitly inform about limb twisting (eg. twisting the upper arm while keeping the lower arm fixed relatively to it would either increase the twisted appearance of the lower arm (eg. if the arm is straight) or lead to a global change of the 3D position of the wrist (eg. if the elbow is bent)). It should be noted that the classification of some other rotations could be integrated as new posecodes, for instance the orientation of the head. We do not claim the designed list of posecodes and super-posecodes to be exhaustive, but the pipeline to be modular and open to new additions. While there will always be room for improvement of the automatic captioning pipeline, pretraining with the captions it produces in the current state already proved highly beneficial.
@@ Appendix C Miscellaneous discussions
![[Uncaptioned image]](extracted/5357568/figures/bio/ginger.jpg)
![[Uncaptioned image]](extracted/5357568/figures/bio/PW.jpg)
![[Uncaptioned image]](extracted/5357568/figures/bio/thomas.jpg)
![[Uncaptioned image]](extracted/5357568/figures/bio/fmoreno.png)
![[Uncaptioned image]](extracted/5357568/figures/bio/greg.jpg)

### v3

@@ A-B List of posecodes
Posecodes statistics. In Figures 16, 17, 18, 19, 20, 21 and 22 we show posecode statistics obtained over PoseScript-A20. Specifically, circle areas represent the proportion of poses satisfying the corresponding posecode categorization for the associated keypoints. We use the black and grey colors to denote categorizations that are ignored in the captioning process. A black circle area means that the corresponding pose configuration is too ambiguous (e.g. when the relative distance between two body parts is close to 0, making the detection of the body parts’ relative position less obvious.). Grey circle areas denote trivial pose configurations (e.g. when a left body part is at the left of the associated right body part: this is the case most of the time). They correspond to posecode categorizations that apply to at least 60% of the poses. In contrast, posecode categorizations that describe less than 6% of the poses are defined as unskippable (i.e., such pose information cannot be randomly discarded during the posecode selection process), and are colored in orange. All other available posecodes categorizations, in blue, are skippable (i.e., such pose information can be randomly discarded during the posecode selection process). Equivalent information for super-posecodes is provided in Table IX.
@@ Appendix B More about the automatic captioning pipeline
Rules were mined based on PoseScript-A20. We further reviewed all mined rules manually, to keep only the most meaningful and dispose of the following:
@@
Rotation-related information. As it uses 3D coordinates as input, the automatic captioning pipeline describes relative body part positioning, but does not inform about limb orientation or limb twist. Yet, this could be easily taken into account by designing new posecodes which computation would rely on joint rotations (instead of joint coordinates). This is mostly interesting and worth describing for end limbs like lower arms (e.g. to tell whether the inside of the lower arm is facing down or facing up, when the wrist is extended forward). Usually, twists of other limbs either induce visual twists of the end limbs connected to them, or result in different positions of other 3D keypoints. As such, in most of the cases, the captioning of these 3D keypoints does implicitly inform about limb twisting. For instance, twisting the upper arm while keeping the lower arm fixed relatively to it would either increase the twisted appearance of the lower arm (e.g. if the arm is straight) or lead to a global change of the 3D position of the wrist (e.g. if the elbow is bent). It should be noted that the classification of some other rotations could be integrated as new posecodes, for instance the orientation of the head. We do not claim the designed list of posecodes and super-posecodes to be exhaustive, but the pipeline to be modular and open to new additions. While there will always be room for improvement of the automatic captioning pipeline, the produced automatic captions are only used for pretraining and already prove to be highly beneficial in their current state.
@@ Appendix C Miscellaneous discussions
Pretraining on PoseScript-A. Since the studied tasks require semantic understanding of the pose, using PoseScript-A for pretraining makes a lot of sense: indeed, it pairs poses with texts, which are semantic by nature. To emphasize the value of this pretraining, we trained a variational auto-encoder (VAE) model on the poses from PoseScript-A, and used its weights to initialize the pose encoder of our retrieval model from section IV, before finetuning it on PoseScript-H (EXP-1). We also experimented finetuning a retrieval model pretrained on PoseScript-A, where only the weights of the pose encoder would be imported (limiting the initialization to the same layers that could be initialized in the previous experiment); EXP-2. These two models can be compared to a model learned from scratch (no pretraining; EXP-0) and a model whose both encoders (the pose’s and text’s) are initialized with the pretrained weights (EXP-3). Results are presented in Table X. It can be observed that any pretraining of the sole pose encoder is useful (+10 points at least on the mRecall, when comparing EXP-1 or EXP-2 to EXP-0), with the VAE-pretraining EXP-1 being better in R@1 and the other “semantic” pretraining EXP-2 being better in R@50. Eventually, initializing both encoder weights with those of a model pretrained on PoseScript-A is the most beneficial (+78%, EXP-3), compared to only initializing the pose encoder (+49%, EXP-1). This highlights the value of PoseScript-A as a set of pose and text pairs for pretraining, and thus validates the contribution of the automatic captioning pipeline.
| EXP | Pretraining | mRecall | pose-to-text | text-to-pose | ||
| 0 | none | 23.0 0.6 | 8.9 0.3 | 34.8 0.5 | 9.3 1.0 | 35.7 0.9 |
| 1 | pose encoder (via VAE) | 34.2 0.2 | 15.5 0.7 | 48.7 0.5 | 15.5 0.5 | 50.2 0.1 |
| 2 | pose encoder | 33.8 1.1 | 14.2 1.2 | 49.1 1.7 | 13.8 0.4 | 50.9 0.8 |
| 3 | pose & text encoders | 40.9 0.1 | 19.8 0.4 | 56.2 0.7 | 19.9 0.6 | 57.9 0.3 |
![[Uncaptioned image]](extracted/5845636/figures/bio/ginger.jpg)
![[Uncaptioned image]](extracted/5845636/figures/bio/PW.jpg)
![[Uncaptioned image]](extracted/5845636/figures/bio/thomas.jpg)
![[Uncaptioned image]](extracted/5845636/figures/bio/fmoreno.png)
![[Uncaptioned image]](extracted/5845636/figures/bio/greg.jpg)